Testing:

H0:p=0.57
H1:p≠0.57

Your sample consists of 138 subjects, with 77 successes. Calculate the test statistic, rounded to 2 decimal places

z=_____
Answer #1

solution

this is two tailed test

= x / n = 77/138=0.5580

P0 = 0.57

1 - P0 = 1-0.57=0.43

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.5580-0.57/ [(057*0.43) / 138]

z= - 0.29

Test statistic = z = -0.29